### v2

@@ 2.2 Integral representations of the density
×$\quad\times I_{\frac{n}{2}-1}\bigg{(}\bigg{|}\frac{\mu_{X}}{\sigma_{X}}-\frac{\mu_{Y}}{\sigma_{Y}}\bigg{|}\frac{\sqrt{{\color[rgb]{0,0,0}nxt}}}{(1-\rho)\sqrt{\sigma_{X}\sigma_{Y}}}\bigg{)}I_{\frac{n}{2}-1}\bigg{(}\bigg{|}\frac{\mu_{X}}{\sigma_{X}}+\frac{\mu_{Y}}{\sigma_{Y}}\bigg{|}\frac{\sqrt{{\color[rgb]{0,0,0}nx(1+t)}}}{(1+\rho)\sqrt{\sigma_{X}\sigma_{Y}}}\bigg{)}\,{\rm d}t,$ (2.12)
@@
$D_{1}(x)$ $=\frac{1}{\sigma_{X}\sigma_{Y}(1-\rho^{2})}\bigg{(}\frac{n}{{\color[rgb]{0,0,0}4}}\bigg{)}^{1-n/2}\bigg{|}\frac{\mu_{X}^{2}}{\sigma_{X}^{2}}-\frac{\mu_{Y}^{2}}{\sigma_{Y}^{2}}\bigg{|}^{1-n/2}\bigg{(}\frac{x}{\sigma_{X}\sigma_{Y}}\bigg{)}^{n/2}$
×$\quad\times\exp\bigg{(}-{\color[rgb]{0,0,0}\frac{n}{2(1-\rho^{2})}\bigg{(}\frac{\mu_{X}^{2}}{\sigma_{X}^{2}}+\frac{\mu_{Y}^{2}}{\sigma_{Y}^{2}}-\frac{2\rho\mu_{X}\mu_{Y}}{\sigma_{X}\sigma_{Y}}\bigg{)}}-\frac{x}{\sigma_{X}\sigma_{Y}(1+\rho)}\bigg{)}\,.$
@@
×$\quad\times I_{\frac{n}{2}-1}\bigg{(}\frac{2|\mu_{X}|}{\sigma_{X}}\frac{\sqrt{{\color[rgb]{0,0,0}nx(1+t)}}}{(1+\rho)\sqrt{\sigma_{X}\sigma_{Y}}}\bigg{)}\,{\rm d}t,$ (2.13)
@@
$D_{2}(x)$ $=\frac{{\color[rgb]{0,0,0}n}^{(2-n)/4}}{\sigma_{X}\sigma_{Y}(1-\rho)^{n/2}(1+\rho)\Gamma(n/2)}\bigg{(}\frac{|\mu_{X}|}{\sigma_{X}}\bigg{)}^{1-n/2}\bigg{(}\frac{x}{\sigma_{X}\sigma_{Y}}\bigg{)}^{(3n-2)/4}$
@@
×$\quad\times I_{\frac{n}{2}-1}\bigg{(}\frac{2|\mu_{X}|}{\sigma_{X}}\frac{\sqrt{{\color[rgb]{0,0,0}nxt}}}{(1-\rho)\sqrt{\sigma_{X}\sigma_{Y}}}\bigg{)}\,{\rm d}t,$ (2.14)
@@
$D_{3}(x)$ $=\frac{{\color[rgb]{0,0,0}n}^{(2-n)/4}}{\sigma_{X}\sigma_{Y}(1+\rho)^{n/2}(1-\rho)\Gamma(n/2)}\bigg{(}\frac{|\mu_{X}|}{\sigma_{X}}\bigg{)}^{1-n/2}\bigg{(}\frac{x}{\sigma_{X}\sigma_{Y}}\bigg{)}^{(3n-2)/4}$
@@
As usual, we prove the result for the case $\sigma_{X}=\sigma_{Y}=1$, with the general case following from a simple rescaling. We will derive formula (2.8), which holds provided $|\mu_{X}|\not=|\mu_{Y}|$ (in the case $\sigma_{X}=\sigma_{Y}=1$). A slight modification of our derivation yields the formulas given in parts 2 and 3 of the theorem under the conditions $\mu_{X}=\mu_{Y}$ and $\mu_{X}=-\mu_{Y}$, respectively (in the case $\sigma_{X}=\sigma_{Y}=1$). We omit the details.
@@
$f_{S_{n}}(x)$ $=C(x)\sum_{k=0}^{\infty}\dfrac{\left(n/{\color[rgb]{0,0,0}8}\right)^{k}}{k!}\bigg{(}\frac{1-\rho}{1+\rho}(\mu_{X}+\mu_{Y})^{2}\bigg{)}^{k}\sum_{j=0}^{k}\frac{1}{\Gamma(n/2+k-j)}\binom{k}{j}\bigg{(}\frac{1+\rho}{1-\rho}\frac{\mu_{X}-\mu_{Y}}{\mu_{X}+\mu_{Y}}\bigg{)}^{2j}$
@@
$f_{S_{n}}(x)$ $=C(x)z^{n-1}\sum_{k=0}^{\infty}\dfrac{\left(nz/{\color[rgb]{0,0,0}8}\right)^{k}}{k!}\bigg{(}\frac{1-\rho}{1+\rho}(\mu_{X}+\mu_{Y})^{2}\bigg{)}^{k}$
@@
Applying the integral representation (A.31) of the confluent hypergeometric function of the second kind, followed by a legitimate changing the order of summation and integration, results in
$f_{S_{n}}(x)$ $=\frac{C(x)z^{n-1}}{\Gamma^{2}(n/2)}\int_{0}^{\infty}{\rm e}^{-zt}t^{n/2-1}(1+t)^{n/2-1}\sum_{k=0}^{\infty}\dfrac{[nz(1+t)]^{k}}{{\color[rgb]{0,0,0}8}^{k}\,k!}\bigg{(}\frac{1-\rho}{1+\rho}(\mu_{X}+\mu_{Y})^{2}\bigg{)}^{k}$
@@
$=\frac{C(x)z^{n-1}}{\Gamma^{2}(n/2)}\int_{0}^{\infty}{\rm e}^{-zt}t^{n/2-1}(1+t)^{n/2-1}\sum_{k=0}^{\infty}\dfrac{[nz(1+t)]^{k}}{{\color[rgb]{0,0,0}8}^{k}\,(n/2)_{k}\,k!}\bigg{(}\frac{1-\rho}{1+\rho}(\mu_{X}+\mu_{Y})^{2}\bigg{)}^{k}$
@@
Now, by virtue of Watson’s formula [31, p. 148, Eq. (2)], viz.
$\sum_{k=0}^{\infty}\frac{(-1)^{k}y^{2k}}{(a)_{k}\,k!}\,{}_{2}F_{1}(-a-k+1,-k;b;x^{2})=\Gamma(a)\,\Gamma(b)\,x^{1-b}\,y^{2-a-b}\,J_{a-1}(2y)\,J_{b-1}(2xy)\,,$ (2.16)
@@
$f_{S_{n}}(x)$ $=\frac{\tilde{D}_{1}(x)}{{\rm i}^{n-2}}\int_{0}^{\infty}(t(t+1))^{(n-2)/4}\exp\bigg{(}-\frac{2xt}{1-\rho^{2}}\bigg{)}\,J_{\frac{n}{2}-1}\bigg{(}\mathrm{i}\sqrt{\frac{1+\rho}{1-\rho}}|\mu_{X}-\mu_{Y}|\sqrt{nzt{\color[rgb]{0,0,0}/2}}\bigg{)}$
×$\quad\times J_{\frac{n}{2}-1}\bigg{(}\mathrm{i}\sqrt{\frac{1-\rho}{1+\rho}}|\mu_{X}+\mu_{Y}|\sqrt{nz(1+t){\color[rgb]{0,0,0}/2}}\bigg{)}\,{\rm d}t,$
where $\tilde{D}_{1}(x)$ is the coefficient $D_{1}(x)$ with $\sigma_{X}=\sigma_{Y}=1$. Finally, using that $J_{\nu}({\rm i}w)={\rm i}^{\nu}\,I_{\nu}(w)$ (which is immediate from the definitions (A.20) and (A.21)) yields the formula (2.8) in the case $\sigma_{X}=\sigma_{Y}=1$, as required.
@@ 2.3 Infinite divisibility
Let $\mu_{X},\mu_{Y}\in\mathbb{R}$, $\sigma_{X},\sigma_{Y}>0$ and $-1<\rho<1$. Then, for any $m\geq 1$, we have the decomposition $Z=_{d}\sum_{i=1}^{m}U_{i,m}$, where $U_{1,m},\ldots,U_{m,m}$ are independent and identically distributed random variables with PDF
$f_{U_{1,m}}(x)$ $=\frac{(1-\rho^{2})^{q/2-1}}{2^{n-1}\sigma_{X}\sigma_{Y}}\exp\bigg{(}-{\color[rgb]{0,0,0}\frac{q}{2(1-\rho^{2})}\bigg{(}\frac{\mu_{X}^{2}}{\sigma_{X}^{2}}+\frac{\mu_{Y}^{2}}{\sigma_{Y}^{2}}-\frac{2\rho\mu_{X}\mu_{Y}}{\sigma_{X}\sigma_{Y}}\bigg{)}}-\frac{|x|}{\sigma_{X}\sigma_{Y}(1+\rho\,\mathrm{sgn}(x))}\bigg{)}$
×$\quad\times\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(q/{\color[rgb]{0,0,0}8})^{k}}{k!\Gamma(q/2+a_{j,k}(x))}\binom{k}{j}\bigg{(}\frac{1+\rho}{1-\rho}\bigg{)}^{j}\bigg{(}\frac{\mu_{X}}{\sigma_{X}}-\frac{\mu_{Y}}{\sigma_{Y}}\bigg{)}^{2j}\bigg{(}\frac{1-\rho}{1+\rho}\bigg{)}^{k-j}$
@@ Appendix A Special functions
$I_{n+\frac{1}{2}}(x)$ $=\frac{1}{\sqrt{2\pi x}}\bigg{\{}\sum_{j=0}^{m}\frac{(-1)^{j}(n+j)!}{j!(n-j)!}\frac{\mathrm{e}^{x}}{(2x)^{j}}+(-1)^{n+1}\sum_{j=0}^{\infty}\frac{(n+j)!}{j!(n-j)!}\frac{\mathrm{e}^{-x}}{(2x)^{j}}\bigg{\}},\quad n\in\mathbb{N}_{0}.$ (A.23)
The functions $J_{\nu}(x)$ and $I_{\nu}(x)$ having the following limiting behaviour as $x\rightarrow 0$,
$J_{\nu}(x)$ $\sim\frac{x^{\nu}}{2^{\nu}\Gamma(\nu+1/2)},\quad\nu>-1,$
$I_{\nu}(x)$ $\sim\frac{x^{\nu}}{2^{\nu}\Gamma(\nu+1/2)},\quad\nu>-1.$ (A.24)
The modified Bessel function of the second kind can be defined, for $\nu\in\mathbb{R}$, by the Laplace–transform type integral [31, p. 183, Eq.(15)] [^5] By Watson, this integral representation of the modified Bessel function of the second kind was probably firstly considered by Poisson in an equivalent form in [25, p. 237].
@@
The following formula corrects the definite integral formula given by [26, p. 325, Eq. 19]:
@@
where $y>0$, $z>0$; $\rho+\sigma>1$; $\delta=\sigma$, $\theta=y$ for $x<0$, and $\delta=\rho$, $\theta=z$ for $x\geq 0$. Here the correction is that the formula of [26] contains an additional factor of $x^{2}$, which should not be present. The following special case of the integral formula (A.37) is given by equation 3.384(9) of [15]:
@@ Acknowledgements
RG is funded in part by EPSRC grant EP/Y008650/1 and EPSRC grant UKRI068. TP was partially supported by the University of Rijeka under the project uniri-iskusni-prirod-23-98.

### v3

@@ 2.2 Integral representations of the density
×$\quad\times I_{\frac{n}{2}-1}\bigg{(}\bigg{|}\frac{\mu_{X}}{\sigma_{X}}-\frac{\mu_{Y}}{\sigma_{Y}}\bigg{|}\frac{\sqrt{nxt}}{(1-\rho)\sqrt{\sigma_{X}\sigma_{Y}}}\bigg{)}I_{\frac{n}{2}-1}\bigg{(}\bigg{|}\frac{\mu_{X}}{\sigma_{X}}+\frac{\mu_{Y}}{\sigma_{Y}}\bigg{|}\frac{\sqrt{nx(1+t)}}{(1+\rho)\sqrt{\sigma_{X}\sigma_{Y}}}\bigg{)}\,{\rm d}t,$ (2.11)
@@
$D_{1}(x)$ $=\frac{1}{\sigma_{X}\sigma_{Y}(1-\rho^{2})}\bigg{(}\frac{n}{4}\bigg{)}^{1-n/2}\bigg{|}\frac{\mu_{X}^{2}}{\sigma_{X}^{2}}-\frac{\mu_{Y}^{2}}{\sigma_{Y}^{2}}\bigg{|}^{1-n/2}\bigg{(}\frac{x}{\sigma_{X}\sigma_{Y}}\bigg{)}^{n/2}$
×$\quad\times\exp\bigg{(}-\frac{n}{2(1-\rho^{2})}\bigg{(}\frac{\mu_{X}^{2}}{\sigma_{X}^{2}}+\frac{\mu_{Y}^{2}}{\sigma_{Y}^{2}}-\frac{2\rho\mu_{X}\mu_{Y}}{\sigma_{X}\sigma_{Y}}\bigg{)}-\frac{x}{\sigma_{X}\sigma_{Y}(1+\rho)}\bigg{)}\,.$
@@
×$\quad\times I_{\frac{n}{2}-1}\bigg{(}\frac{2|\mu_{X}|}{\sigma_{X}}\frac{\sqrt{nx(1+t)}}{(1+\rho)\sqrt{\sigma_{X}\sigma_{Y}}}\bigg{)}\,{\rm d}t,$ (2.12)
@@
$D_{2}(x)$ $=\frac{n^{(2-n)/4}}{\sigma_{X}\sigma_{Y}(1-\rho)^{n/2}(1+\rho)\Gamma(n/2)}\bigg{(}\frac{|\mu_{X}|}{\sigma_{X}}\bigg{)}^{1-n/2}\bigg{(}\frac{x}{\sigma_{X}\sigma_{Y}}\bigg{)}^{(3n-2)/4}$
@@
×$\quad\times I_{\frac{n}{2}-1}\bigg{(}\frac{2|\mu_{X}|}{\sigma_{X}}\frac{\sqrt{nxt}}{(1-\rho)\sqrt{\sigma_{X}\sigma_{Y}}}\bigg{)}\,{\rm d}t,$ (2.13)
@@
$D_{3}(x)$ $=\frac{n^{(2-n)/4}}{\sigma_{X}\sigma_{Y}(1+\rho)^{n/2}(1-\rho)\Gamma(n/2)}\bigg{(}\frac{|\mu_{X}|}{\sigma_{X}}\bigg{)}^{1-n/2}\bigg{(}\frac{x}{\sigma_{X}\sigma_{Y}}\bigg{)}^{(3n-2)/4}$
@@
As usual, we prove the result for the case $\sigma_{X}=\sigma_{Y}=1$. We will derive formula (2.8), which holds provided $|\mu_{X}|\not=|\mu_{Y}|$ (in the case $\sigma_{X}=\sigma_{Y}=1$). A slight modification of our derivation yields the formulas given in parts 2 and 3 of the theorem under the conditions $\mu_{X}=\mu_{Y}$ and $\mu_{X}=-\mu_{Y}$, respectively We omit the details.
@@
$f_{S_{n}}(x)$ $=C(x)\sum_{k=0}^{\infty}\dfrac{\left(n/8\right)^{k}}{k!}\bigg{(}\frac{1-\rho}{1+\rho}(\mu_{X}+\mu_{Y})^{2}\bigg{)}^{k}\sum_{j=0}^{k}\frac{1}{\Gamma(n/2+k-j)}\binom{k}{j}\bigg{(}\frac{1+\rho}{1-\rho}\frac{\mu_{X}-\mu_{Y}}{\mu_{X}+\mu_{Y}}\bigg{)}^{2j}$
@@
$f_{S_{n}}(x)$ $=C(x)z^{n-1}\sum_{k=0}^{\infty}\dfrac{\left(nz/8\right)^{k}}{k!}\bigg{(}\frac{1-\rho}{1+\rho}(\mu_{X}+\mu_{Y})^{2}\bigg{)}^{k}$
@@
Applying the integral representation (A.29) followed by changing the order of summation and integration results in
$f_{S_{n}}(x)$ $=\frac{C(x)z^{n-1}}{\Gamma^{2}(n/2)}\int_{0}^{\infty}{\rm e}^{-zt}t^{n/2-1}(1+t)^{n/2-1}\sum_{k=0}^{\infty}\dfrac{[nz(1+t)]^{k}}{8^{k}\,k!}\bigg{(}\frac{1-\rho}{1+\rho}(\mu_{X}+\mu_{Y})^{2}\bigg{)}^{k}$
@@
$=\frac{C(x)z^{n-1}}{\Gamma^{2}(n/2)}\int_{0}^{\infty}{\rm e}^{-zt}t^{n/2-1}(1+t)^{n/2-1}\sum_{k=0}^{\infty}\dfrac{[nz(1+t)]^{k}}{8^{k}\,(n/2)_{k}\,k!}\bigg{(}\frac{1-\rho}{1+\rho}(\mu_{X}+\mu_{Y})^{2}\bigg{)}^{k}$
@@
Now, by virtue of Watson’s formula [32, p. 148, Eq. (2)], viz.
$\sum_{k=0}^{\infty}\frac{(-1)^{k}y^{2k}}{(a)_{k}\,k!}\,{}_{2}F_{1}(-a-k+1,-k;b;x^{2})=\Gamma(a)\,\Gamma(b)\,x^{1-b}\,y^{2-a-b}\,J_{a-1}(2y)\,J_{b-1}(2xy)\,,$
@@
$f_{S_{n}}(x)$ $=\frac{\tilde{D}_{1}(x)}{{\rm i}^{n-2}}\int_{0}^{\infty}(t(t+1))^{(n-2)/4}\exp\bigg{(}-\frac{2xt}{1-\rho^{2}}\bigg{)}\,J_{\frac{n}{2}-1}\bigg{(}\mathrm{i}\sqrt{\frac{1+\rho}{1-\rho}}|\mu_{X}-\mu_{Y}|\sqrt{nzt/2}\bigg{)}$
×$\quad\times J_{\frac{n}{2}-1}\bigg{(}\mathrm{i}\sqrt{\frac{1-\rho}{1+\rho}}|\mu_{X}+\mu_{Y}|\sqrt{nz(1+t)/2}\bigg{)}\,{\rm d}t,$
where $\tilde{D}_{1}(x)$ is the coefficient $D_{1}(x)$ with $\sigma_{X}=\sigma_{Y}=1$, and we recall that $z=2x/(1-\rho^{2})$. Finally, using that $J_{\nu}({\rm i}w)={\rm i}^{\nu}\,I_{\nu}(w)$ (which is immediate from the definitions (A.18) and (A.19)) yields the formula (2.8) in the case $\sigma_{X}=\sigma_{Y}=1$, as required.
@@ 2.3 Infinite divisibility
For any $m\geq 1$, we have the decomposition $Z=_{d}\sum_{i=1}^{m}U_{i,m}$, where $U_{1,m},\ldots,U_{m,m}$ are independent and identically distributed random variables with PDF
$f_{U_{1,m}}(x)$ $=\frac{(1-\rho^{2})^{q/2-1}}{2^{n-1}\sigma_{X}\sigma_{Y}}\exp\bigg{(}-\frac{q}{2(1-\rho^{2})}\bigg{(}\frac{\mu_{X}^{2}}{\sigma_{X}^{2}}+\frac{\mu_{Y}^{2}}{\sigma_{Y}^{2}}-\frac{2\rho\mu_{X}\mu_{Y}}{\sigma_{X}\sigma_{Y}}\bigg{)}-\frac{|x|}{\sigma_{X}\sigma_{Y}(1+\rho\,\mathrm{sgn}(x))}\bigg{)}$
×$\quad\times\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(q/8)^{k}}{k!\Gamma(q/2+a_{j,k}(x))}\binom{k}{j}\bigg{(}\frac{1+\rho}{1-\rho}\bigg{)}^{j}\bigg{(}\frac{\mu_{X}}{\sigma_{X}}-\frac{\mu_{Y}}{\sigma_{Y}}\bigg{)}^{2j}\bigg{(}\frac{1-\rho}{1+\rho}\bigg{)}^{k-j}$
@@ Appendix A Special functions
$I_{n+\frac{1}{2}}(x)$ $=\frac{1}{\sqrt{2\pi x}}\bigg{\{}\sum_{j=0}^{n}\frac{(-1)^{j}(n+j)!}{j!(n-j)!}\frac{\mathrm{e}^{x}}{(2x)^{j}}+(-1)^{n+1}\sum_{j=0}^{n}\frac{(n+j)!}{j!(n-j)!}\frac{\mathrm{e}^{-x}}{(2x)^{j}}\bigg{\}}.$ (A.21)
The function $I_{\nu}(x)$ has the following limiting behaviour as $x\rightarrow 0$,
$I_{\nu}(x)$ $\sim\frac{x^{\nu}}{2^{\nu}\Gamma(\nu+1)},\quad\nu>-1.$ (A.22)
The modified Bessel function of the second kind can be defined, for $\nu\in\mathbb{R}$, by the Laplace–transform type integral [32, p. 183, Eq.(15)] [^5] By Watson, this integral representation of the modified Bessel function of the second kind was probably firstly considered by Poisson in an equivalent form in [26, p. 237].
@@
The following formula corrects the definite integral formula given by [27, p. 325, Eq. 19]:
@@
where $y>0$, $z>0$; $\rho+\sigma>1$; $\delta=\sigma$, $\theta=y$ for $x<0$, and $\delta=\rho$, $\theta=z$ for $x\geq 0$. The formula (A.35) is also valid in the case $\rho=\sigma=1/2$, which can be seen from Lemma 3.1 of [24] and an application of the formula (A.31). The correction we make is that the formula of [27] contains an additional factor of $x^{2}$, which should not be present. The formula (A.35) can alternatively be derived from equation 3.384(9) of [16] together with an application of formula (A.34); the formula of [16] is expressed in terms of the Whittaker function rather than the confluent hypergeometric function of the second kind. The following special case of the integral formula (A.35) is given by equation 3.384(9) of [16]:
@@ Acknowledgements
RG is funded in part by EPSRC grant EP/Y008650/1 and EPSRC grant UKRI068. TP was partially supported by the University of Rijeka under the project uniri-iskusni-prirod-23-98. We would like to thank the reviewer for carefully reading our paper and for their helpful comments and suggestions.